# Equivalent class of Emergent Single Weyl Fermion in 3d Topological States: gapless superconductors and superfluids Vs chiral fermions

### v1

@@ III Naturalness of single Weyl fermion in gapless superconductors or superfluids
Physically, this suggests naturalness of single Weyl fermion in gapless charge $U(1)$ symmetry breaking states such as gapless superconductors or superfluids. They shall form an equivalent class. Later in Sect.VI, we provide an explicit structure of the equivalent class of Hamiltonians.
@@
Here we show an alternative proof of the celebrated Nielsen-Ninomyia no-go theorem [43, 42]. The proof presents a fundamentally different approach to the problem than the original one [42] and the other ones that we could find in the literature [44],[45]. Despite requiring some knowledge on differential geometry, this proof is actually quite natural. We present the idea of the proof here (with a $2\times 2$ Hamiltonian) and the complete mathematical proof, as well as some refinements, will be reported in the Appendix A.
@@
We place ourselves in the 6-dimensional manifold $\mathbb{T}^{3}\times\mathfrak{su}(2)$, in which we define the 3-dimensional submanifolds $S:=\{(p,H(p))\mid p\in T^{3}\}$ and $\mathbb{T}^{3}\times\{0\}$, which are automatically oriented.
@@ IV Fermion Lattice Model Analysis I
A simple lattice model of a time reversal symmetric tQCP with protection symmetry $G_{p}=Z_{2}^{T}$, the time reversal symmetry group. And we restrict ourselves to a generic tQCP where the change of topologies is minimum. The change takes a fundamental value consistent with the protecting symmetry $G_{p}$ which in this case is $\delta N^{f}_{w}=2$ (the superscript refers to the fundamental value of changes). The specific lattice model describes a tQCP in the DIII class topological superconductors. The IR limit of such a tQCP has been studied quite extensively in Ref.[28, 33] and its lattice model in a connection to a 3d boundary of a 4d topological state was also presented[41].
@@
Applying a strong magnetic field to a fully gaped SPT again in the DIII class with protecting symmetry $G_{p}=Z_{2}^{T}$, i.e. a topological superconductor or superfluid that break the charge $U(1)$ symmetry. Quantum phase transitions into these gapless superconducting nodal points in Ref.[27, 28] and a lattice model was presented in a unpublished thesis[32].
@@
The last three models involve either a generic tQCP or fine-tuned multi-critical tQCPs which are further subject to applications of time reversal symmetry breaking actions. It can be thought as a hydride of Path a) and Path b). In Type c), we study lattice models of tQCPs with the change of topologies taking values of integer multiples of the fundamental values (i.e. $\delta N^{f}_{w}=2$), i.e. we work with tQCP models with $\delta N_{W}=4,8$ in additional to a generic tQCP with $\delta N_{w}=\delta N^{f}_{w}=2$. A lattice model that was originally proposed a few months ago in [53] belongs to one of Type c models in our classification with $\delta N_{w}=8$.
We will focus of the UV completion of the emergent infrared single Weyl cone physics in gapless superfluids and corresponding conserved charge operators in lattice models in the momentum space of Torus-three, $\mathcal{T}^{3}$.
@@
Here $N^{f}_{D}(G_{p})$ is the number of fundamental degrees of freedom when the change of topologies takes a fundamental value. $\delta N^{f}_{w}=\delta N^{f}_{w}(G_{p})$ again is the fundamental value of the change of topologies defined by the protecting symmetry $G_{p}$$\delta N_{w}$ is the actual change of topologies at a tQCP which interests us.
@@
As we have forcast before, and will be further seen below, in both Path b) and Path c), the time reversal symmetry has to be broken in the lattice models, in addtional to the charge $U(1)$ spontaneously symmetry breaking to make the construction feasible. Only in Path a), the T-symmetry can be fully preserved.
In this section, we will first focus on Path a) and Path b), and illustrate the ideas via two simplest lattice models respectively. The infrared properties of gapless superfluids in these models were studied before and are known. And as stated before, here we will focus on their UV completion and the UV completed symmetry group.
@@ IV.1 Lattice Model I:
$Q(\mathbf{k})=$ $\sin(k_{x})\sigma^{z}-\sin k_{y}\sigma^{x}-\sin k_{z}\tau^{y}\otimes\sigma^{y}$ (46)
@@
in the infrared limit of ${\bf k}\rightarrow 0$. This is an emergent chiral charge for a tQCP[28]. As predicted for any chiral charge of a single Weyl fermion, the charge goes to 0 at $\mathbf{k}=(\pi,\pi,\pi)$. It is possible to find the real space charge by expanding the last term in a Fourier series, but for the sake of brevity we do not show the result here. Instead, we can have the following alternative charge:
$Q(\mathbf{k})=$ $\sum_{i}\sin^{2}(k_{i}/2)\times\left(\sin(k_{x})\sigma^{z}-\sin k_{y}\sigma^{x}-\sin k_{z}\tau^{y}\otimes\sigma^{y}\right)$ (47)
@@
but has a simpler charge with second nearest neighbor couplings:
$\mathcal{Q}=\sum_{i}$ $[\left(3\chi_{i+\hat{x}}-\frac{1}{2}(\sum_{b}\chi_{i+\hat{x}+b}+\chi_{i+\hat{x}-b})\right)i\sigma^{z}\chi_{i}$ (48)
$-\left(3\chi_{i+\hat{y}}-\frac{1}{2}(\sum_{b}\chi_{i+\hat{y}+b}+\chi_{i+\hat{y}-b})\right)i\sigma^{x}\chi_{i}$
$-\left(3\chi_{i+\hat{z}}-\frac{1}{2}(\sum_{b}\chi_{i+\hat{z}+b}+\chi_{i+\hat{z}-b})\right)i\tau^{y}\sigma^{y}\chi_{i}$
$-\left(3\chi_{i}-\chi_{i+2\hat{x}}-\chi_{i+2\hat{y}}-\chi_{i+2\hat{z}}\right)\tau^{x}\sigma^{y}\chi_{i}]+h.c.$
@@ IV.2 Lattice model II:
We add a magnetic field to a gaped SPT with Time-Reversal symmetry (corresponding to the previous model with $\epsilon\neq 0$), therefore producing a superconductor or superfluid that breaks the charge U(1) symmetry. We choose a particular representation of the $\Gamma$ matrices of section II.3, so that the Hamiltonian looks like a lattice completion of a p-wave superconductor Hamiltonian (which presents explicit charge U(1) symmetry breaking). The model studied is the following:
@@
$\mathcal{H}=\mu\sum_{x,y,z}c_{x,y,z}^{\dagger}c_{x,y,z}-\frac{1}{2}\sum_{x,y,z}\left\{c_{x,y,z}^{\dagger}c_{x+1,y,z}+c_{x,y,z}^{\dagger}c_{x,y+1,z}+c_{x,y,z}^{\dagger}c_{x,y,z+1}+\text{h.c.}\right\}+\sum_{x,y,z}c_{x,y,z}^{\dagger}\left(\vec{B}\cdot\vec{\sigma}\right)c_{x,y,z}$ (50)
$-\frac{1}{2}\sum_{x,y,z}\left\{c_{x,y,z}(\sigma^{z})c_{x+1,y,z}+ic_{x,y,z}c_{x,y+1,z}-c_{x,y,z}(\sigma^{x})c_{x,y,z+1}+\text{h.c.}\right\}$
@@
$\mathcal{Q}=\sum_{i}$ $[\left(\frac{1}{2}(\sum_{b}\chi_{i+\hat{z}+b}+\chi_{i+\hat{z}-b})-\mu\chi_{i+\hat{z}}\right)iI\otimes\sigma^{z}\chi_{i}$ (52)
$+\frac{1}{2}\left(\chi_{i}-\chi_{i+2\hat{z}}\right)\tau^{x}\otimes\sigma^{y}\chi_{i}]+h.c.$
@@ V.3 Model V: a tQCP with $\delta N_{W}=8$ subject to T-symmetry breaking actions
These touching points are related to quantum critical points of Lifschitz type and can disturb the IR dynamics as discussed in Ref.[27]. The low energy dynamics in the specific construction therefore are characterized by a mix of two different dynamic critical exponents $z=1,2$. The low energy sector further contains gapless Lifchitz fermions, in addition to a single Weyl fermion.
Nevertheless, one can remove these gapless Lifschitz points by increasing the TRB field,
@@
VI Organizing lattice models in a family: An equivalent class of single Weyl fermion models
We now further show that the infrared limit for the models in the different discussions can be organized into a family of Hamiltonians which transforms in a $3\otimes 3$ dimensional linear representation of $\text{Spin}(4)$ group [28] where one of the $SU(2)$ subgroups can be identified as a subgroup of the emergent Lorentz $SO(3,1)$ group.
The symmetry group for real fermions is $\text{Spin(4) = SU(2)}\timesSU$(2), so as we are considering theories of real fermions, we should consider the action of Spin(4) on the real fermion Hamiltonians studied. the spin(4) algebra is generated by two su(2) algebras that commute with each other, we choose a particular realization of their generators:
@@ VI Organizing lattice models in a family: An equivalent class of single Weyl fermion models
In the following, we will show these actions of $Spin(4)$ explicitly in a tensor product representation. We will also present two dual copies of such representations and illustrate that all chiral lattice fermion models introduced in Sect.II,IV,V can be organized into these two representations and form an equivalent class.
@@ VI.2 The mass operators in lattice models
They are simply the generators of two $su(2)$ subgroups of $spin(4)$ defined in Eq.67,68. They themselves form a $3\oplus 3$ adjoint representations of $Spin(4)$ group spanning a six dimensional space.
@@ VI.3 The time reversal symmetry breaking fields
Therefore, the whole lattice model structure can be simply encoded in the substructure of $H$, the effective Hamiltonian defined in Eq.79,80. In conclusion, the lattice chiral fermions with a single weyl cone can therefore form an equivalent class defined by Eq.79,80,82,85,86,88,83.
@@
However, all the chiral fermion lattice models with a single Weyl fermion, after all symmetry constraints along with the Clifford algebra are imposed can be uniquely encoded in two dual copies of the much smaller 9-dimensional $(1,1)$ representations defined in Eq.79,80. So effectively, the $15\oplus 15$ dimensional manifold can be projected into a $9\oplus 9$ dimensional sub-manifold where chiral fermion lattice models with single Weyl fermions discussed here all belong to.
@@
where the bars refer to the $\tau-\sigma$ or charge-spin dual transformation. In a sense, its infrared substructures contains unique information of the whole family of lattice models. And all models form an equivalent class either defined by a tQCP in the DIII class topological superconductors or by its dual, a nodal point phase, i.e. they are isomorphic.
@@ VII Generalized UV completed symmetries and symmetry charge operators
Symmetry charges or charge-operators have distinct structures and span distinguishable spaces in models that are either non-degenerate along path b), path c), i.e. the models in Eq.49,Eq.53,Eq.57,Eq.64 or degenerate along path a), i.e. the model in Eq.45. We show that the space of conserved charges is larger in the latter.
@@ VII.1 Charges for $\cal T$}-symmetry breaking Hamiltonians
For the first case, consider for example Hamiltonians in Eq.49, 53,57,64. At nearly all points or any generic point in the Brillouin zone, their eigenvalues are non-degenerate, this indicates that any charge operators that commute with $H$ must obey:
@@
$\{Q_{I}({\bf k}),Q_{II}({\bf k});{\bf k}\in\mathcal{T}^{3}\}$ (92)
defined at different points of $\mathcal{T}^{3}$ are $\mathbf{k}$- dispersive and also contain the null operator. And these charges defined above form rays spanning $\mathbb{R}\otimes\mathbb{R}$ (see below) rather than in a compact manifold.

### v2

@@ III Naturalness of single Weyl fermion in gapless superconductors or superfluids
Physically, this suggests naturalness of single Weyl fermion in gapless charge $U(1)$ symmetry breaking states such as gapless superconductors or superfluids. They shall form an equivalence class. Later in Sect.VI, we provide an explicit structure of the equivalence class of Hamiltonians.
@@
Here we show an alternative proof of the celebrated Nielsen-Ninomyia no-go theorem [39, 38]. The proof presents a fundamentally different approach to the problem than the original one [38] and the other ones that we could find in the literature [19],[29]. The intuitive topological proof given by Nielsen and Ninomiya [39] uses intersection theory as well, but with a projective line and a 3-torus embedded in a four dimensional space. Our approach is more systematic and has the advantage of showing the handedness of the Weyl points directly from the formalism. We believe it is potentially generalizable to other dimensions. Despite requiring some knowledge on differential geometry, this proof is actually quite natural. We present the idea of the proof here (with a $2\times 2$ Hamiltonian) and the complete mathematical proof, as well as some refinements, will be reported in the Appendix A.
@@
We place ourselves in the 6-dimensional manifold $\mathbb{T}^{3}\times\mathfrak{su}(2)$, in which we define the 3-dimensional submanifolds $S:=\{(p,H(p))\mid p\in\mathbb{T}^{3}\}$ and $\mathbb{T}^{3}\times\{0\}$, which are automatically oriented.
@@ IV Fermion Lattice Model Analysis I
A simple lattice model of a time reversal symmetric tQCP with protection symmetry $G_{p}=Z_{2}^{T}$, the time reversal symmetry group. And we restrict ourselves to a generic tQCP where the change of topologies is minimum. The change of band toplogies here physically is related to the recontruction of surface majorana cones and dictates the gapless fermions at tQCPs (see 43) [^1] Here We label the majorana band topologies here via $N_{w}=2n,n=0,1,2,$.... In literature, the band topologies sometimes are also labeled by $\nu$ which is related to $N_{w}$ via $N_{w}=2\nu$.
The change takes a fundamental value consistent with the protecting symmetry $G_{p}$ which in this case is $\delta N^{f}_{w}=2$ (the superscript refers to the fundamental value of changes). The specific lattice model describes a tQCP in the DIII class topological superconductors. The IR limit of such a tQCP has been studied quite extensively in Ref.[62, 61] and its lattice model in a connection to a 3d boundary of a 4d topological state was also presented[63].
@@
Applying a strong magnetic field to a fully gapped SPT again in the DIII class with protecting symmetry $G_{p}=Z_{2}^{T}$, i.e. a topological superconductor or superfluid that break the charge $U(1)$ symmetry. Quantum phase transitions into these gapless superconducting nodal points in Ref.[58, 62] and a lattice model was presented in a unpublished thesis[28].
@@
The last three models involve either a generic tQCP or fine-tuned multi-critical tQCPs which are further subject to applications of time reversal symmetry breaking actions. It can be thought as a hybrid of Path a) and Path b). In Type c), we study lattice models of tQCPs with the change of topologies taking values of integer multiples of the fundamental values (i.e. $\delta N^{f}_{w}=2$), i.e. we work with tQCP models with $\delta N_{W}=4,8$ in additional to a generic tQCP with $\delta N_{w}=\delta N^{f}_{w}=2$. A lattice model that was originally proposed a few months ago in [21] belongs to one of Type c models in our classification with $\delta N_{w}=8$.
We will focus on the UV completion of the emergent infrared single Weyl cone physics in gapless superfluids and corresponding conserved charge operators in lattice models in the 3-torus momentum space, $\mathbb{T}^{3}$.
@@
Here $N^{f}_{D}(G_{p})$ is the number of fundamental degrees of freedom when the change of topologies takes a fundamental value. $\delta N^{f}_{w}=\delta N^{f}_{w}(G_{p})$ again is the fundamental value of the change of topologies defined by the protecting symmetry $G_{p}$. $\delta N_{w}$ is the actual change of topologies at a tQCP which interests us.
@@
As we have predicted before, and will be further seen below, in both Path b) and Path c), the time reversal symmetry has to be broken in the lattice models, in addition to the charge $U(1)$ spontaneously breaking the symmetry to make the construction feasible. Only in Path a), the T-symmetry can be fully preserved.
In this section, we will first focus on Path a) and Path b), and illustrate the ideas via two simplest lattice models, respectively. The infrared properties of gapless superfluids in these models have been previously studied and are known. And, as stated before, here we will focus on their UV completion and the UV completed symmetry group.
@@ IV.1 Lattice Model I:
$Q_{1}(\mathbf{k})=$ $\sin(k_{x})\sigma^{z}-\sin k_{y}\sigma^{x}-\sin k_{z}\tau^{y}\otimes\sigma^{y}$ (46)
@@
in the infrared limit of ${\bf k}\rightarrow 0$. This is an emergent chiral charge for a tQCP[62]. As predicted for any chiral charge of a single Weyl fermion, the charge goes to 0 at $\mathbf{k}=(\pi,\pi,\pi)$. It is possible to find the real space charge by expanding the last term in a Fourier series:
$c_{\bf r}=\int_{BZ}\frac{\sum_{i}\sin^{2}(k_{i})}{\sum_{i}\sin^{2}(k_{i}/2)}e^{-i\bf r\cdot\bf k}d\bf k$
Then the real space charge writes:
$\mathcal{Q}_{1}=\sum_{\bf i}[$ $\chi_{{\bf i}+\hat{x}}i\sigma^{z}\chi_{\bf i}-\chi_{{\bf i}+\hat{y}}i\sigma^{x}\chi_{\bf i}-\chi_{{\bf i}+\hat{z}}i\tau^{y}\otimes\sigma^{y}\chi_{\bf i}$ (47)
$-\sum_{\bf r}c_{\bf r}\chi_{{\bf i}+{\bf r}}\tau^{x}\otimes\sigma^{y}\chi_{\bf i}]+h.c.$
with $\bf r$} going through the whole lattice. We use $\hat{x},\hat{y},\hat{z}$ to denote the neighboring sites along the $x,y,z$-directions respectively.
This charge has a complicated real-space spatial structure because of the k-dependence in Eq.46. Alternatively, we can have the following charge:
$Q_{2}(\mathbf{k})=$ $\sum_{i}\sin^{2}(k_{i}/2)\times\left(\sin(k_{x})\sigma^{z}-\sin k_{y}\sigma^{x}-\sin k_{z}\tau^{y}\otimes\sigma^{y}\right)$ (48)
@@
which has a simpler real space expression with second nearest neighbor couplings:
$\mathcal{Q}_{2}=\sum_{\bf i}$ $[\left(3\chi_{{\bf i}+\hat{x}}-\frac{1}{2}(\sum_{\hat{b}}\chi_{{\bf i}+\hat{x}+\hat{b}}+\chi_{{\bf i}+\hat{x}-\hat{b}})\right)i\sigma^{z}\chi_{\bf i}$ (49)
$-\left(3\chi_{{\bf i}+\hat{y}}-\frac{1}{2}(\sum_{\hat{b}}\chi_{{\bf i}+\hat{y}+\hat{b}}+\chi_{{\bf i}+\hat{y}-\hat{b}})\right)i\sigma^{x}\chi_{\bf i}$
$-\left(3\chi_{{\bf i}+\hat{z}}-\frac{1}{2}(\sum_{\hat{b}}\chi_{{\bf i}+\hat{z}+\hat{b}}+\chi_{{\bf i}+\hat{z}-\hat{b}})\right)i\tau^{y}\sigma^{y}\chi_{\bf i}$
$-\left(3\chi_{\bf i}-\chi_{{\bf i}+2\hat{x}}-\chi_{{\bf i}+2\hat{y}}-\chi_{{\bf i}+2\hat{z}}\right)\tau^{x}\sigma^{y}\chi_{\bf i}]+h.c.$
with $\hat{b}$ only going through the neighboring sites.
@@ IV.2 Lattice model II:
We add a magnetic field to a gapped SPT with Time-Reversal symmetry (corresponding to the previous model with $\epsilon\neq 0$), therefore producing a superconductor or superfluid that breaks the charge U(1) symmetry. We choose a particular representation of the $\Gamma$ matrices of section II.3, so that the Hamiltonian looks like a lattice completion of a p-wave superconductor Hamiltonian (which presents explicit charge U(1) symmetry breaking). The model studied is the following:
@@
$\mathcal{H}=\mu\sum_{\bf i}c_{\bf i}^{\dagger}c_{\bf i}-\frac{1}{2}\sum_{\bf i}\left\{c_{\bf i}^{\dagger}c_{{\bf i}+\hat{x}}+c_{\bf i}^{\dagger}c_{{\bf i}+\hat{y}}+c_{\bf i}^{\dagger}c_{{\bf i}+\hat{z}}+\text{h.c.}\right\}+\sum_{\bf i}c_{\bf i}^{\dagger}\left(\vec{B}\cdot\vec{\sigma}\right)c_{\bf i}$ (51)
$-\frac{1}{2}\sum_{\bf i}\left\{c_{\bf i}(\sigma^{z})c_{{\bf i}+\hat{x}}+ic_{\bf i}c_{{\bf i}+\hat{y}}-c_{\bf i}(\sigma^{x})c_{{\bf i}+\hat{z}}+\text{h.c.}\right\}$
@@
$\mathcal{Q}=\sum_{\bf i}$ $[\left(\frac{1}{2}(\sum_{\hat{b}}\chi_{{\bf i}+\hat{z}+{\hat{b}}}+\chi_{{\bf i}+\hat{z}-{\hat{b}}})-\mu\chi_{{\bf i}+\hat{z}}\right)iI\otimes\sigma^{z}\chi_{\bf i}$ (53)
$+\frac{1}{2}\left(\chi_{\bf i}-\chi_{{\bf i}+2\hat{z}}\right)\tau^{x}\otimes\sigma^{y}\chi_{\bf i}]+h.c.$
@@ V.3 Model V: a tQCP with $\delta N_{W}=8$ subject to T-symmetry breaking actions
These touching points are related to quantum critical points of Lifshitz type and can disturb the IR dynamics as discussed in Ref.[58]. The low energy dynamics in the specific construction therefore are characterized by a mix of two different dynamic critical exponents $z=1,2$. The low energy sector further contains gapless Lifshitz fermions, in addition to a single Weyl fermion.
Nevertheless, one can remove these gapless Lifshitz points by increasing the TRB field,
@@
VI Organizing lattice models in a family: An equivalence class of single Weyl fermion models
We now further show that the infrared limit for the models in the different discussions can be organized into a family of Hamiltonians which transform in a $3\otimes 3$ dimensional linear representation of $\text{Spin}(4)$ group [62] where one of the $SU(2)$ subgroups can be identified as a subgroup of the emergent Lorentz $SO(3,1)$ group.
The unitary group acting on these real fermions here is $\text{Spin(4) = SU(2)}\timesSU$(2), so as we are considering theories of real fermions, we should consider the action of Spin(4) on the real fermion Hamiltonians studied. the spin(4) algebra is generated by two su(2) algebras that commute with each other, we choose a particular realization of their generators:
@@ VI Organizing lattice models in a family: An equivalence class of single Weyl fermion models
In the following, we will show these actions of $Spin(4)$ explicitly in a tensor product representation. We will also present two dual copies of such representations and illustrate that all chiral lattice fermion models introduced in Sect.II,IV,V can be organized into these two representations and form an equivalence class.
@@ VI.2 The mass operators in lattice models
They are simply the generators of two $su(2)$ subgroups of $spin(4)$ defined in Eq.68,69. They themselves form a $3\oplus 3$ adjoint representations of $Spin(4)$ group spanning a six dimensional space.
@@ VI.3 The time reversal symmetry breaking fields
Therefore, the whole lattice model structure can be simply encoded in the substructure of $H$, the effective Hamiltonian defined in Eq.80,81. In conclusion, the lattice chiral fermions with a single weyl cone can therefore form an equivalence class defined by Eq.80,81,83,86,87,89,84.
@@
However, all the chiral fermion lattice models with a single Weyl fermion, after all symmetry constraints along with the Clifford algebra are imposed can be uniquely encoded in two dual copies of the much smaller 9-dimensional $(1,1)$ representations defined in Eq.80,81. So effectively, the $15\oplus 15$ dimensional manifold can be projected into a $9\oplus 9$ dimensional sub-manifold where chiral fermion lattice models with single Weyl fermions discussed here all belong to.
@@
where the bars refer to the $\tau-\sigma$ or charge-spin dual transformation. In a sense, its infrared substructures contains unique information of the whole family of lattice models. And all models form an equivalence class either defined by a tQCP in the DIII class topological superconductors or by its dual, a nodal point phase, i.e. they are isomorphic.
@@ VII Generalized UV completed symmetries and symmetry charge operators
Symmetry charges or charge-operators have distinct structures and span distinct spaces in models that are either non-degenerate along path b), path c), i.e. the models in Eq.50,Eq.54,Eq.58,Eq.65 or degenerate along path a), i.e. the model in Eq.45. We show that the space of conserved charges is larger in the latter.
@@ VII.1 Charges for $\cal T$}-symmetry breaking Hamiltonians
For the first case, consider for example Hamiltonians in Eq.50, 54,58,65. At nearly all points or any generic point in the Brillouin zone, their eigenvalues are non-degenerate, this indicates that any charge operators that commute with $H$ must obey:
@@
$\{Q_{I}({\bf k}),Q_{II}({\bf k});{\bf k}\in\mathbb{T}^{3}\}$ (93)
defined at different points of $\mathbb{T}^{3}$ are $\mathbf{k}$- dispersive and also contain the null operator. And these charges defined above form rays spanning $\mathbb{R}\otimes\mathbb{R}$ (see below) rather than in a compact manifold.